# Martingale Approach to Gambler’s Ruin Problem for Correlated Random Walks

### v1

@@
Abstract The gambler’s ruin problem for correlated random walks, both with and without delays, is addressed by Optional Stopping Theorem for martingales.
@@ 2 Correlated Random Walk
Let $S_{n}=X_{1}+\cdots+X_{n}$ denote a correlated random walk or CRW, and let ${\mathcal{F}}_{n}=\sigma(X_{1},\dots,X_{n})$, the $\sigma$-field generated by $\{X_{n}\}_{n\geq 1}$. For any integers $A,B>0$, let us consider the following stopping time (w.r.t to the filtration $\{{\mathcal{F}}_{n}\}_{n\geq 1}$):
@@
One important technical observation is that ${\mathbf{E}}(\tau)<\infty$. It is a well-known fact, so we will skip the proof. Then the “ruin” probabilities are: $\alpha={\mathbf{P}}(S_{\tau}=A)$ and $\beta={\mathbf{P}}(S_{\tau}=-B)$.
When $p=1-q$, $\{X_{n}\}_{n\geq 1}$ are independent identically distributed (iid) random variables. The gambler’s ruin problem can be solved with the help of various techniques. The most popular ones are first-step analysis and the martingale approach. A review of both methods can be found in Steele (2001). In particular, the following martingales are useful. For a symmetric random walk ($p=q=1/2$), they are $S_{n}$ and $S_{n}^{2}-n$. For asymmetric random walks, they are $(q/p)^{S_{n}}$ and $S_{n}-(p-q)n$. Then by a standard application of the Optional Stopping Theorem (OST) one can show that, for example, in the symmetric case $\alpha=B/(A+B)$ and ${\mathbf{E}}(\tau)=AB$.
The general Markovian case was first treated in Mohan (1955) with the help of the first-step analysis. To the best of our knowledge, nobody has used a martingale approach to address the gambler’s ruin problem for correlated random walks. So, our first task is to re-derive the formulas from Mohan (1955) with the help of the Optional Stopping Theorem. For that, we need to construct an additive martingale (similar to $S_{n}-(p-q)n$ in the iid case).
@@ 2.1 Ruin Probability for Symmetric CRW
$0={\mathbf{E}}(M_{1})={\mathbf{E}}(M_{\tau})=(2-2p){\mathbf{E}}(S_{\tau})+(2p-1){\mathbf{E}}(X_{\tau})$
or
@@
    if $p\to 1$, then $\alpha\to 1/2$ (everything is decided by the first flip)
@@ 2.2 Expected Waiting Time until Ruin for Symmetric CRW
The iid case tells us that we should try to find a “quadratic” martingale. This suggests the following candidate:
@@
$\left\{\begin{array}[]{l}2(2p-1)+a(2p-1)=a\\ 1+a-bn=-b(n-1)\end{array}\right.$
@@ 2.3 Ruin Probability for Asymmetric CRW
The iid case together with our first martingale construction suggests an “exponential” martingale
@@
$\left\{\begin{array}[]{l}p\lambda+(1-p)\lambda^{-1-2a}=1\\ (1-q)\lambda^{1+2a}+q\lambda^{-1}=1\end{array}\right.$
Because of the iid martingale $(q/p)^{S_{n}}$, it is natural to try $\lambda=q/p$. And it works, $\lambda=q/p$ and
@@
${\mathbf{E}}(M_{1})=$ $\left(\frac{q}{p}\right)^{1+a}\frac{1-q}{2-p-q}+\left(\frac{q}{p}\right)^{-1-a}\frac{1-p}{2-p-q}=\frac{\sqrt{(1-q)(1-p)}}{2-p-q}\left[\sqrt{\frac{q}{p}}+\sqrt{\frac{p}{q}}\right]$
@@
If $p=1-q$, $p\neq 1/2$ , then formula (6) simplifies to a well-known expression for the asymmetric iid case (for example, (Steele, 2001)).
@@ 2.4 Expected Waiting Time until Ruin for Asymmetric CRW
Since we already know $\alpha$ (and $\beta=1-\alpha$) and ${\mathbf{E}}(M_{1})={\mathbf{E}}(X_{1})=(p-q)/(2-p-q)$, the formula for ${\mathbf{E}}(\tau)$ immediately follows.
@@ 3 Correlated Random Walk with Delays
As we mentioned in the introduction, accounting for the Markovian dependence is not enough to solve our two-pattern problem. The payments are quite often equal to 0. Random walks associated with such random variables are called “random walks with delays”. The gambler’s ruin problem with delays for the iid case was solved in Gut (2013) with the help of martingales and the OST. Here we extend these results to CRW with delays.
@@ 3.1 Ruin Probability for Symmetric CRW with Delays
We will keep the same notation for random walks, ruin probabilities, filtration and stopping time. We call this random walk symmetric, because states 1 and -1 enter into the picture in a symmetric way, they are interchangeable.
@@
So, a natural candidate for another “useful” martingale is
@@
    if $12p+r=$ 1 or $p=1-p-r$ (almost the iid case because of an additional symmetry), then $S_{n}$ itself a martingale, and $\alpha=B/(A+B)$
@@ 3.2 Expected Waiting Time until Ruin for Symmetric CRW with Delays
Finding ${\mathbf{E}}(\tau)$ is more challenging, because the delays play a significant role. At this point, we have some experience in building “useful” martingales, and this experience suggests the following candidate:
@@
Note that because of the restrictions on $p$, $q$, and $r$, there is no division by zero. The last step is to calculate ${\mathbf{E}}(M_{1})$. Since ${\mathbf{E}}(X_{1}^{2})=(1-q)/(1+r-q)$, we get
@@
    if $12p+r=$ 1 and $r=q$ (the iid symmetric random walk with delays), then ${\mathbf{E}}(\tau)=AB/(1-q)$, a formula from Gut (2013).
@@ 3.3 Ruin Probability for General CRW with Delays
The candidate for a martingale in the general case is given by
@@
If one can find $\lambda\neq 1$, $a$ and $b$ that make $M_{n}$ a martingale, then the OST gives us

### v2

@@
Abstract The gambler’s ruin problem for correlated random walks (CRW), both with and without delays, is addressed using the Optional Stopping Theorem for martingales. We derive closed-form expressions for the ruin probabilities and the expected game duration for CRW with increments $\{1,-1\}$ and for symmetric CRW with increments $\{1,0,-1\}$ (CRW with delays). Additionally, a martingale technique is developed for general CRW with delays. The gambler’s ruin probability for a game involving bets on two arbitrary patterns is also examined.
@@
1 Introduction
A correlated random walk is a sequence of partial sums of random variables whose increments form a Markov chain. Thus, it is a natural generalization of the standard random walk. The CRW is also a special case of a Markov random walk (for example, Grama et al. (2018) and Fuh (2004)). As it was noticed by Gut and Stadtmüller (2021), a symmetric CRW with increments $\{1,-1\}$ can also be viewed a special case of an elephant random walk when the elephant only remembers the very last step. In this paper we investigate a one-dimensional CRW; a $d$-dimensional CRW is called the Gillis-Domb-Fisher correlated random walk (see Chen and Renshaw (1992) and Gillis (1955)).
CRW has numerous applications in statistical physics, finance, ecology and other areas of research. Here are just a few examples. Allaart (2004) employs CRW to maximize the expected return of a commodity seller. Gardner et al. (2022) uses it to model short-term directional persistence in animal movements. Applications to quantum walks in connection with quantum computing are used as motivation in Konno (2009). Our interest in random walks with delays (that is, with $\{1,0,-1\}$ increments) is prompted by applications to sequential detection in sensor networks with communication constraints (see Guerriero et al. (2010)). In this work, we consider a scenario where a sensor takes a measurement and, due to battery limitations, proceeds as follows: if the measurement is too high, above a certain threshold, it sends a 1 to the fusion center; if it is too low, it sends a -1. When the measurement is within an acceptable range, the sensor does nothing, which is interpreted as a passive response, or 0.
The gambler’s ruin problem for CRW was first investigated in Mohan (1955) with the help of first-step analysis. The case of a two-state $k$-order Markov chain was studied in Mukherjea and Steele (1987) by the difference equations method. However, as they noted, handling a large number of difference equations is difficult even for a small $k$. A numerical method (via $n$-step transition probabilities) for finding the ruin probabilities and the expected duration of the game is given by Lal and Bhat (1989). The gambler’s ruin problem with delays for the independent increment case was solved in Gut (2013) with the help of martingales.
The paper proceeds as follows. In Section 2 we describe a game that is used as an illustration of a Markovian gambler’s ruin problem with delays. In Section 3 we derive closed-form expressions for both the ruin probabilities and the expected time until ruin for the correlated random walk with increments $\{1,-1\}$. Section 4 includes closed-form expressions for the ruin probabilities and the expected time until ruin for symmetric CRW with delays and a martingale treatment of the general case. Section 5 provides a solution for the gambler’s ruin problem in the case of the two arbitrary pattern game. In the final section we briefly discuss some open questions.
@@ 2 A Two-Pattern Problem
Let us introduce a simple game that naturally leads us to the gambler’s ruin problem for correlated random walks with delays. This game will be used to demonstrate the flexibility and limitations of our methodology. The complete solution to this gambler’s ruin problem will be provided in Section 5, along with a solution for the general case.
@@
We must acknowledge that there is an element of luck in the sequence of payments in this particular game forming a Markov chain with three states $\{1,0,-1\}$. Usually, this is not the case. One can easily verify that if we play the same game with HH and TT, the payments do not form a Markov chain. However, as we demonstrate later, finding the ruin probability for a game with two arbitrary patterns is closely related to the gambler’s ruin problem for the CRW.
@@ 3 Correlated Random Walk
Let $S_{n}=X_{1}+\cdots+X_{n}$, $n\geq 1$ denote a correlated random walk, and let ${\mathcal{F}}_{n}=\sigma(X_{1},\dots,X_{n})$, the $\sigma$-field generated by $\{X_{n}\}_{n\geq 1}$. For any integers $A,B>0$, let us consider the following stopping time (w.r.t to the filtration $\{{\mathcal{F}}_{n}\}_{n\geq 1}$):
@@
One important technical observation is that ${\mathbf{E}}(\tau)<\infty$. It is a well-known fact, so we will skip the proof. Then the ruin probabilities are: $\alpha={\mathbf{P}}(S_{\tau}=A)$ and $\beta={\mathbf{P}}(S_{\tau}=-B)$.
When $p=1-q$, $\{X_{n}\}_{n\geq 1}$ are independent identically distributed (i.i.d.) random variables. The gambler’s ruin problem can be solved with the help of various techniques. The most popular ones are first-step analysis and the martingale approach. A review of both methods can be found in Steele (2001). In particular, the following martingales are useful. For a symmetric random walk ($p=q=1/2$), they are $S_{n}$ and $S_{n}^{2}-n$. For asymmetric random walks, they are $(q/p)^{S_{n}}$ and $S_{n}-(p-q)n$. Then by a standard application of the Optional Stopping Theorem (OST) one can show that, for example, in the symmetric case $\alpha=B/(A+B)$ and ${\mathbf{E}}(\tau)=AB$.
The general Markovian case was first treated in Mohan (1955) with the help of the first-step analysis. To the best of our knowledge, nobody has used a martingale approach to address the gambler’s ruin problem for correlated random walks. So, our first task is to re-derive the formulas from Mohan (1955) with the help of the Optional Stopping Theorem. For that, we need to construct an additive martingale (similar to $S_{n}-(p-q)n$ in the i.i.d. case).
@@ 3.1 Ruin Probability for Symmetric CRW
$0={\mathbf{E}}(M_{1})={\mathbf{E}}(M_{\tau})=(2-2p){\mathbf{E}}(S_{\tau})+(2p-1){\mathbf{E}}(X_{\tau}).$
Since $X_{\tau}=1$ on the set $\{S_{\tau}=A\}$ , and $X_{\tau}=-1$ on $\{S_{\tau}=-B\}$ we have
@@
    if $p\to 1$, then $\alpha\to 1/2$ (everything is decided by the first flip),
  4. 4.
    if $B\to\infty$, then $\alpha\to 1$.
As mentioned above, adjusting to an arbitrary initial distribution $(\pi_{1},1-\pi_{1})$ is straightforward. The only change is the value of ${\mathbf{E}}(M_{1})$. In this case, ${\mathbf{E}}(M_{1})=2\pi_{1}-1$. Consequently, we have
$\alpha=\frac{B-1+\frac{\pi_{1}}{1-p}}{A+B+\frac{2p-1}{1-p}}.$ (3)
@@ 3.2 Expected Waiting Time until Ruin for Symmetric CRW
The i.i.d. case tells us that we should try to find a quadratic martingale. This suggests the following candidate:
@@
$\left\{\begin{array}[]{l}2(2p-1)+a(2p-1)=a\\ 1+a-bn=-b(n-1)\end{array}\right..$
@@ 3.3 Ruin Probability for Asymmetric CRW
The i.i.d. case together with our first martingale construction suggests an exponential martingale
@@
$\left\{\begin{array}[]{l}p\lambda+(1-p)\lambda^{-1-2a}=1\\ (1-q)\lambda^{1+2a}+q\lambda^{-1}=1\end{array}\right..$
Because of the i.i.d. martingale $(q/p)^{S_{n}}$, it is natural to try $\lambda=q/p$. And it works, $\lambda=q/p$ and
@@
${\mathbf{E}}(M_{1})=$ $\left(\frac{q}{p}\right)^{1+a}\frac{1-q}{2-p-q}+\left(\frac{q}{p}\right)^{-1-a}\frac{1-p}{2-p-q}=\frac{\sqrt{(1-q)(1-p)}}{2-p-q}\left[\sqrt{\frac{q}{p}}+\sqrt{\frac{p}{q}}\right].$
@@
If $p=1-q$, $p\neq 1/2$ , then formula (7) simplifies to a well-known expression for the asymmetric i.i.d. case (for example, (Steele, 2001)):
$\alpha=\frac{\left(\frac{q}{p}\right)^{-B}-1}{\left(\frac{q}{p}\right)^{-B}-\left(\frac{q}{p}\right)^{A}}.$
@@ 3.4 Expected Waiting Time until Ruin for Asymmetric CRW
Since we already know $\alpha$ (and $\beta=1-\alpha$) and ${\mathbf{E}}(M_{1})={\mathbf{E}}(X_{1})=(p-q)/(2-p-q)$, the formula for ${\mathbf{E}}(\tau)$ immediately follows:
${\mathbf{E}}(\tau)=\frac{2-p-q}{p-q}(A\alpha-B\beta)+\frac{p+q-1}{p-q}(\alpha-\beta)+\frac{1-p-q}{2-p-q}.$
@@ 4 Correlated Random Walk with Delays
As we mentioned in the introduction, accounting for the Markovian dependence is not enough to solve our two-pattern problem. The payments are quite often equal to 0. Random walks associated with such random variables are called random walks with delays. The gambler’s ruin problem with delays for the i.i.d. case can be solved using standard techniques of martingales and the OST, see Gut (2013). Here we extend these results to CRW with delays. We start with a symmetric CRW which can be viewed as a natural generalization of the i.i.d. symmetric random walk with delays.
@@ 4.1 Ruin Probability for Symmetric CRW with Delays
We will keep the same notation for random walks, ruin probabilities, filtration and stopping time. We call this random walk symmetric, because states 1 and -1 enter into the picture in a symmetric way, they are interchangeable. More specifically, for states 1 and -1, the probabilities of staying in the same state and the probabilities of switching to state 0 are the same. Additionally, the probabilities of switching from 0 to 1 and -1 are equal, too.
@@
So, a natural candidate for another useful martingale is
@@
    if $12p+r=$ 1 or $p=1-p-r$ (almost the i.i.d. case because of an additional symmetry), then $S_{n}$ itself a martingale, and $\alpha=B/(A+B)$
@@ 4.2 Expected Waiting Time until Ruin for Symmetric CRW with Delays
Finding ${\mathbf{E}}(\tau)$ is more challenging, because the delays play a significant role. At this point, we have some experience in building useful martingales, and this experience suggests the following candidate:
@@
Note that because of the restrictions on $p$, $q$, and $r$, there is no division by zero. The last step is to calculate ${\mathbf{E}}(M_{1})$. Since we start from the stationary distribution, ${\mathbf{E}}(X_{1}^{2})=(1-q)/(1+r-q)$, and, as a result, we get
@@
    if $12p+r=$ 1 and $r=q$ (the i.i.d. symmetric random walk with delays), then ${\mathbf{E}}(\tau)=AB/(1-q)$, a formula from Gut (2013).
@@ 4.3 Ruin Probability for General CRW with Delays
The candidate for a martingale in the general case, as before, takes an exponential form:
@@
If one can find $\lambda>0$, $\lambda\neq 1$, $a$ and $b$ that make $M_{n}$ a martingale, then the OST gives us
@@
$=$ ${\mathbf{E}}(Y_{n}|X_{n-1}=1)\frac{X_{n-1}^{2}+X_{n-1}}{2}+{\mathbf{E}}(Y_{n}|X_{n-1}=-1)\frac{X_{n-1}^{2}-X_{n-1}}{2}$
$+{\mathbf{E}}(Y_{n}|X_{n-1}=0)(1-X_{n-1}^{2})$
$=$ $\mathbin{\hphantom{+}}\Big{[}{\mathbf{E}}(Y_{n}|X_{n-1}=1)+{\mathbf{E}}(Y_{n}|X_{n-1}=-1)-2{\mathbf{E}}(Y_{n}|X_{n-1}=0)\Big{]}\frac{X_{n-1}^{2}}{2}$
$+\Big{[}{\mathbf{E}}(Y_{n}|X_{n-1}=1)-{\mathbf{E}}(Y_{n}|X_{n-1}=-1)\Big{]}\frac{X_{n-1}}{2}$
$+\Big{[}{\mathbf{E}}(Y_{n}|X_{n-1}=0)\Big{]}.$
Taking into account that
${\mathbf{E}}(Y_{n}|X_{n-1}=1)=p\lambda+p_{0}\lambda^{-a-b}+(1-p-p_{0})\lambda^{-2b-1},$
${\mathbf{E}}(Y_{n}|X_{n-1}=0)=r_{1}\lambda^{a+b+1}+r+(1-r-r_{1})\lambda^{a-b-1},$
${\mathbf{E}}(Y_{n}|X_{n-1}=-1)=(1-q-q_{0})\lambda^{2b+1}+q_{0}\lambda^{-a+b}+q\lambda^{-1},$
we finally get the following system of non-linear equations
$\left\{\begin{array}[]{lr}p\lambda+p_{0}\lambda^{-a-b}+(1-p-p_{0})\lambda^{-2b-1}&=1\\ r_{1}\lambda^{a+b+1}+r+(1-r-r_{1})\lambda^{a-b-1}&=1\\ (1-q-q_{0})\lambda^{2b+1}+q_{0}\lambda^{-a+b}+q\lambda^{-1}&=1\\ \end{array}\right..$ (12)
Example 1.
Consider a three-state Markov chain with state set $\{1,0,-1\}$, transition matrix
$\begin{pmatrix}1/2&1/4&1/4\\ 1/3&1/3&1/3\\ 1/8&1/8&3/4\\ \end{pmatrix},$
and the initial distribution $(0,1,0)$. Then one can show that $\lambda=13/10$, $a=-\log_{13/10}\sqrt{81/65}$, and $b=\log_{13/10}\sqrt{20/13}$ solve system (12). Also in this case ${\mathbf{E}}(M_{1})=1$ Plugging these values into formula (11) we get
$\alpha=\frac{(18/13)(13/10)^{B}-1}{(20/13)(13/10)^{A+B}-1}.$
@@
One practical recommendation is to do a re-parametrization, and use $x=\lambda$, $y=\lambda^{a}$, and $z=\lambda^{b}$ instead of the original $\lambda$, $a$ and $b$. Then Mathematica easily finds solutions of the system.
This method might not work for some Markov chains, if their transition probability matrix exhibits a certain symmetry. Recall that in the i.i.d. case, $(q/p)^{S_{n}}$ is useless, if $p=q$. In fact, our two-pattern problem is a case of this nature.
4.4 Expected Waiting Time until Ruin for General CRW with Delays
Using the same technique as before one can show that
${\mathbf{E}}(X_{n}|{\mathcal{F}}_{n-1})=l_{2}X_{n-1}^{2}+l_{1}X_{n-1}+1_{0},$
where
$l_{2}=p-q+(p_{0}-q_{0})/2+1-r-2r_{1},\quad\quad l_{1}=p+q+(p_{0}+q_{0})/2-1,\quad\quad l_{0}=2r_{1}+r-1,$
and
${\mathbf{E}}(X_{n}^{2}|{\mathcal{F}}_{n-1})=k_{2}X_{n-1}^{2}+k_{1}X_{n-1}+k_{0},$
where
$k_{2}=r-(p_{0}+q_{0})/2,\quad\quad k_{1}=(q_{0}-p_{0})/2,\quad\quad k_{0}=1-r.$
Remember that for the asymmetric i.i.d. case a simple martingale $S_{n}-cn$ works. Therefore, it makes sense to try a linear martingale
$M_{n}=S_{n}+aX_{n}^{2}+bX_{n}-cn.$
Using our expressions for ${\mathbf{E}}(X_{n}|{\mathcal{F}}_{n-1})$ and ${\mathbf{E}}(X_{n}^{2}|{\mathcal{F}}_{n-1})$ we can show that if $a$, $b$, and $c$ satisfy
$\left\{\begin{array}[]{lr}ak_{2}+(1+b)l_{2}&=a\\ ak_{1}+(1+b)l_{1}&=b\\ ak_{0}+(1+b)l_{0}&=c\\ \end{array}\right.,$ (13)
then $M_{n}$ is a martingale. Moreover, if $c\neq 0$, then we get a formula the expected time until ruin:
${\mathbf{E}}(\tau)=\frac{1}{c}\big{[}(A+a+b)\alpha+(-B+a-b)(1-\alpha)-{\mathbf{E}}(M_{1})\big{]},$
where
${\mathbf{E}}(M_{1})=\pi_{1}(1+a+b)+(1-\pi_{1}-\pi_{0})(-1+a-b)-c.$
If we consider the Markov chain in Example 1, then our method will lead to the following martingale:
$M_{n}=S_{n}-\frac{2}{5}X_{n}^{2}+\frac{37}{45}X_{n}+\frac{4}{15}n.$
5 Gambler’s Ruin Probability for Two-Pattern Game
In this section, we first provide a complete solution of the HH vs TH gambler’s ruin problem. Then, we describe a martingale method that can be used to find the ruin probabilities for an arbitrary two-pattern game. Finally, we will use the method to re-derive the ruin probability for the HH vs TH game.
5.1 HH vs TH Gambler’s Ruin Problem
Let $\{\xi_{n}\}_{n\geq 1}$ be a sequence of independent Bernoulli random variables with a probability of success 1/2. Then for $n\geq 2$ our payments are given by
$X_{n}=\left\{\begin{array}[]{rl}1,&\mbox{ if }\xi_{n-1}=1,\xi_{n}=1,\\ 0,&\mbox{ if }\xi_{n}=0,\\ -1,&\mbox{ if }\xi_{n-1}=0,\xi_{n}=1.\end{array}\right.$
In the general case of two patterns of length 2 (for instance, HH and TT), the associated sequence of payments typically does not form a Markov chain. But for HH and TH, $\{X_{n}\}_{n\geq 2}$ is a first-order Markov chain with the transition matrix
$\begin{pmatrix}1/2&1/2&0\\ 0&1/2&1/2\\ 1/2&1/2&0\\ \end{pmatrix},$
and the initial distribution $(1/4,1/2,1/4)$. This Markov chain is not symmetric in the sense of Section 4.1, but it does have a very special structure. As a result, one can easily verify that for this transition matrix, there is no solution of system (12) such that $\lambda\neq 1$.
However, we can construct a linear one. Note that here $S_{n}=X_{2}+\cdots+X_{n}$. Since
${\mathbf{E}}(X_{n}|{\mathcal{F}}_{n-1})=X_{n-1}^{2}-\frac{1}{2},$
and
${\mathbf{E}}(X_{n}^{2}|{\mathcal{F}}_{n-1})=\frac{1}{2},$
using our approach (or by solving linear system (13)) one can verify that $M_{n}=S_{n}+X_{n}^{2},\quad n\geq 2$ is a martingale. Applying the OST to this martingale gives us
$\alpha=\frac{B-1/2}{A+B}.$
A simple check in this case is when $A=B=1$. The formula gives 1/4. It makes sense, because the only way to get to 1 is to have HH right away.
Since the linear martingale does not contain an $n$ term, it cannot be used to compute ${\mathbf{E}}(\tau)$. But then we can try to find a quadratic martingale similar to the one used for the symmetric CRW with delays. Indeed, one can show that
$M_{n}=S_{n}^{2}+2S_{n}X_{n}^{2}+2X_{n}^{2}-n/2$
is a martingale. Applying the OST to this martingale gives us that
${\mathbf{E}}(\tau)=2AB+B-A+1.$
5.2 Gambler’s Ruin Probability for Arbitrary Two-Pattern Game
As mentioned above, for an arbitrary two-pattern game the sequence of payments typically is not a Markov chain. For example, suppose that we play this game with HH and TT. As before, let $\{X_{n}\}_{n\geq 2}$ be the sequence of payments, where we get 1 for HH, -1 for TT, and 0, otherwise. Then for any $n\geq 2$, we have
${\mathbf{P}}(X_{2n+1}=1|X_{2n}=0,X_{2n-1}=0,\dots,X_{3}=0,X_{2}=1)=\frac{1}{2},$
but
${\mathbf{P}}(X_{2n+1}=1|X_{2n}=0,X_{2n-1}=0,\dots,X_{3}=0,X_{2}=-1)=0.$
However, we can still find the ruin probability using CRW. For the rest of the section, let $\{\xi_{n}\}_{n\geq 1}$ be an i.i.d. sequence of letters from a finite alphabet. A sequences of letters is called a pattern. A pattern that has no letters is called a void pattern. Consider two patterns
$P=a_{1}a_{2}\dots a_{k},\quad\quad\mbox{and}\quad\quad Q=b_{1}b_{2}\dots b_{m},$
where $k,m\geq 1$. Assume that $P$ and $Q$ are not connected subsequences of each other. Our game is the same as before: we get \$1 for every (overlapping) occurrence of $P$ and we lose a dollar if we see $Q$.
Let $\{X_{n}\}_{n\geq 1}$ be the sequence of only 1 or -1 payments (that is, 0s are ignored). Then $\{X_{n}\}_{n\geq 1}$ is a two-state Markov chain whose initial distribution and transition probabilities can be computed with the help of the theory developed for the occurrence of patterns. More specifically, let
$P^{*}=\left\{\begin{array}[]{rl}a_{2}\dots a_{k},&k\geq 2,\\ \mbox{void pattern},&k=1,\end{array}\right.$
and
$Q^{*}=\left\{\begin{array}[]{rl}b_{2}\dots b_{m},&m\geq 2,\\ \mbox{void pattern},&m=1.\end{array}\right.$
Then the initial probability
$\pi_{1}={\mathbf{P}}(\mbox{pattern $P$ will occur before $Q$ in the sequence$\{\xi_{n}\}_{n\geq 1}$}).$
The transition probability of going from state 1 to state 1 is given by
$p={\mathbf{P}}(\mbox{pattern $P$ will occur before $Q$ in the sequence $\{\xi_{n}\}_{n\geq 1}$ given that it starts from $P^{*}$}),$
and the transition probability of going from state -1 to state -1 is given by
$q={\mathbf{P}}(\mbox{pattern $Q$ will occur before $P$ in the sequence $\{\xi_{n}\}_{n\geq 1}$ given that it starts from $Q^{*}$}).$
All three probabilities can be found using, for example, Theorem 3.1 in Li (1980). To find each probability, it is required to solve a linear system with three equations. Then, we simply apply the appropriate formula from Section 3.
[…]
$\alpha=\frac{B-1/2}{A+B}.$
6 Concluding Remarks
The expected time until the gambler’s ruin in a game with two arbitrary patterns remains an open question. The method based on the occurrence of patterns ignores zero-payments, making it unclear how it can be applied to this problem. Additionally, our technique relies on the fact that indicators of events like $X_{n}=1$ can be expressed in terms of $X_{n}$. Consequently, it is likely unsuitable for Markov chains with more than three states. The distribution of the waiting time until ruin for CRW with delays is an open question as well. Finally, the gambler’s ruin problem for elephant walks is also unresolved. A comprehensive review of the results on elephant random walks can be found in the recent work by Gut and Stadtmüller (2023).